Problem 22-39 (LO. 8, 9) At the beginning of the tax year, Lizzie holds a $10,000 stock basis as the sole shareholder of Spike, Inc., an S corporation. During the year, Spike reports the following: Net taxable income from sales $25,000 Net short-term capital loss (18,000) Cash distribution to Lizzie, 12/31 40,000 If an amount is zero, enter "0". 57,000 X. a. Determine Lizzie's stock basis at the end of the year. $ b. Of Lizzie's $40,000 cash distribution, $ is tax-free and $ is capital gain a c. The entity's short-term capital loss is not deductible by Lizzie this year.
Problem 22-28 (LO. 6) On January 1, 2018, Kinney, Inc., an S corporation, reports $4,000 of accumulated E & P and a balance of $10,000 in AAA. Kinney has two shareholders, Erin and Frank, each of whom owns 500 shares of Kinney's stock. Kinney's nonseparately stated ordinary income for the year is $5,000 Kinney distributes $6,000 to each shareholder on July 1, and it distributes another $3,000 to each shareholder on December 21 How are the shareholders taxed on the distributions? Ignore the 20% QBI deduction Round any division to five decimal places. If required, round final answers to the nearest dollar. Erin and Frank each report $ 10,000 X dividend income for the July 1 distribution and each for the December 21 distribution. Assuming that the shareholders 5,000 have sufficient basis in their stock, both Erin and Frank each receive a tax-free distribution from AAA 7,500
Answer #1

22-39

Part A

Lizzie’s stock basis at the end of year = $0

Beginning stock basis

10000

Plus: Net operating income

25000

Basis subtotal

35000

Minus: Cash distribution

(40000)

Basis subtotal

0

Short-term capital loss

(18000), suspended

Ending stock basis

0

Part B

Of Lizzie’s $40000 cash distributions, $35000 is tax-free and $5000 is a capital gain.

22-28

Erin and Frank each report $1000 dividend income on July 1 distribution and $500 each for the December 21 distribution.

Total AAA = 10000+5000 = 15000

Total distribution = 18000

Therefore, Erin and Frank’s taxed distribution (dividend income) = $3000 (18000-15000)

Portion of AAA allocated to Erin and Frank in July 1 = 12000*15000/18000 = $10000

Portion of AAA allocated to Erin and Frank in December 21 = 6000*15000/18000 = $5000

Therefore ratio is 2:1

In the same ratio, Erin and Frank’s dividend income is distributed

July 1 = 3000*2/3 = $2000 ($1000 each)

December 21 = 3000*1/3 = $1000 ($500 each)
Answer #2

answered by: anonymous